In the laboratory you dissolve 15.1 g of zinc acetate in a volumetric flask and add water to a total volume of 250 . mL.

What is the molarity of the solution?

_________ M
1
Expert's answer
2019-03-14T04:26:48-0400

Zinc acetate - Zn(CH3CO2)2


"C_m= \frac{m}{M \cdot V}"

Cm - molarity or molar concentration of the solution, M

m - mass of the solute, g

M - molar mass of the solute, g/mol

V - volume of the solution, l


"M(Zn(CH_3CO_2)_2) \approx 1 \cdot 65.38 + 2 \cdot (1 \cdot 12.01 + 3 \cdot 1 + 1 \cdot 12.01 + 2 \cdot 16)=183.42 \; g/mol"

"C_m= \frac{15.1}{183.42 \cdot 0.25} \approx 0.33 \; M"


Answer: 0.33 M.